What is the Ionic Strength of a solution that is 0.8M in HCl and 0.05M of H3PO4


1
Expert's answer
2022-04-19T03:11:53-0400

"I=0.5 \sum C_iz_i^2=0.5(1x0.8+1x0.8+ (-3)^2x0.05+1x3x0.05)=1.1"